A medical supply company must appoint a​ president, chief executive officer​ (CEO), chief operating officer​ (COO), and chief financial officer​ (CFO). It must also appoint a planning committee with three different members. There are 14 qualified​ candidates, and officers can also serve on the committee. Complete parts​ (a) through​ (c) below. a. How many different ways can the officers be​ appointed? There are 24024 different ways to appoint the officers. ​(Simplify your​ answer.) b. How many different ways can the committee be​ appointed? There are nothing different ways to appoint the committee. ​(Simplify your​ answer.) c. What is the probability of randomly selecting the committee members and getting the three youngest of the qualified​ candidates? ​P(getting the three youngest of the qualified ​candidates)equals nothing ​(Type an integer or a simplified​ fraction.)
Answer #1

a)

number of ways =14P3 =14*13*12=2184

b)

number of ways committee be​ appointed =14C3 =364

c)

probability of randomly selecting the committee members and getting the three youngest of the qualified​ candidates

=1/364